OpenStudy (anonymous):

Solve the triangle. A = 32°, a = 19, b = 14

OpenStudy (anonymous):

first step is to draw a reasonable triangle

OpenStudy (anonymous):

@$$ angle, side, side, use law of cosines

OpenStudy (anonymous):

hmmm i would use the law of sines, because it is easier

OpenStudy (anonymous):

B = 23°, C = 125°, c ≈ 17.6 Cannot be solved B = 23°, C = 125°, c ≈ 29.4 B = 23°, C = 145°, c ≈ 23.5

OpenStudy (anonymous):



OpenStudy (anonymous):

notice that in the drawing, i made a 32 degree angle look somewhat like 32, maybe not exactly but somewhat also it is important to note that \(a=19\) is longer than \(b=14\)

OpenStudy (anonymous):

yeah sines dodges the ambiguous case

OpenStudy (anonymous):

the via the law of sines you know \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\]so \[\frac{\sin(32)}{19}=\frac{\sin(B)}{14}\] making \[\sin(B)=\frac{14\sin(32)}{19}\]

OpenStudy (anonymous):

you cannot use the law of cosines for this one for the law of sines you need to know 3 out of the 4 numbers in \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\] which you do

OpenStudy (anonymous):

also it is the law of sines that may have an "ambiguous case" not the law of cosines in any case once we have \[\sin(B)=\frac{14\sin(32)}{19}\] we get \(B\) via \[B=\sin^{-1}\left(\frac{14\sin(32)}{19}\right)\]

OpenStudy (anonymous):

i get \(B=22.98\) http://www.wolframalpha.com/input/?i=arcsin%2814sin%2832%29%2F19%29

OpenStudy (anonymous):

which i guess your answers are calling \(23\)

OpenStudy (anonymous):

then you can easily find \(C\) since the angles have to add up to \(180\) and one is \(32\) the other \(23\)